OpenStudy (anonymous):

Need help please i will fan ANY answer and i will medal best answer!!! If y=22, find M

OpenStudy (anonymous):

@misty1212 @welshfella

OpenStudy (anonymous):

@BeastBoy17

OpenStudy (beastboy17):

ok I got a website I use that might help brb

OpenStudy (anonymous):

ok

OpenStudy (beastboy17):

https://mathway.com/ this may help

OpenStudy (anonymous):

ok thanks

OpenStudy (beastboy17):

let me know if it works

OpenStudy (anonymous):

i need help on how to do it though

OpenStudy (beastboy17):

ok hold up whats the question

OpenStudy (anonymous):

If y=22, find M<C.

OpenStudy (beastboy17):

is it a multiple choice or you have to write it ?

OpenStudy (anonymous):

A.44 B.68 C.136 D.60

OpenStudy (beastboy17):

I have someone who could help u understand better

OpenStudy (beastboy17):

@ganeshie8

OpenStudy (beastboy17):

which do you think could be the answer

OpenStudy (beastboy17):

btw

OpenStudy (anonymous):

44???

OpenStudy (mle):

angle A = 22 + 22 = 44 then to find Angle C 180 - 44 = 2*(m<C) solve

OpenStudy (mle):

please ask in the mathematics section next time

OpenStudy (beastboy17):

thanks for the help mLe

OpenStudy (anonymous):

i am still confused can you do it step by step

OpenStudy (mle):

properties of an isosceles triangle the two angles at the top are equal to each other angle A = y + ? = y + y = 22 + 22 get that?

OpenStudy (anonymous):

ok yeah thanks

OpenStudy (mle):

then angles B and C are also equal to each other because it is an isosceles triangle A + B + C = 180 (because it is a triangle) B = C so A + C + C = 180 2C = 180 - A A = (22+22) as we found out before

OpenStudy (mle):

2 C = 180 - 44

OpenStudy (anonymous):

i get 272 though

OpenStudy (mle):

divide 2 from both sides to cancel a multiplied 2

OpenStudy (mle):

if you have something you want to move, to the opposite of what it is currently doing 2 C is currently multiplying to move it, we should divide

OpenStudy (anonymous):

\[2\times44=180-44???\]

OpenStudy (anonymous):

or \[2x=180-44\]

OpenStudy (mle):

we are looking for C only A = 44 2 x C = 180 - 44 C = (180 - 44) / 2 yes

OpenStudy (anonymous):

so 2x=180-44

OpenStudy (mle):

yes

OpenStudy (anonymous):

c=68??

OpenStudy (anonymous):

thanks